# Student Member, IEEE

### v1

@@ Abstract
Received signal strength (RSS) information has seldom been incorporated in the direct position determination (DPD) method of passive radio emitter localization to date. Further, the common use of directional emitters modulates the RSS such that omnidirectional assumptions would dramatically decrease accuracy. This paper introduces a new DPD approach utilizing an RSS-enhanced adaptive beamforming method demonstrating on par or better than state-of-the-art performance at very low SNR for omnidirectional emitters. The technique is then applied to directional emitters taking the imposed RSS modulation into account using a beampattern library, significantly improving localization region confidence as compared to omnidirectional assumption approaches. This is the first approach to date in the open literature for localizing directional emitters.

### v2

@@ Abstract
The localization of RF emitters has attracted significant attention particularly within the domain of electronic warfare. Most localization methods found in open literature are based on omnidirectional emitters. Directional emitters significantly modulate received signal strength (RSS) resulting in degraded performance for localization techniques not modeling this behavior. This paper introduces a direct position determination (DPD) approach utilizing RSS information and adaptive beamforming to localize emitters at very low signal-to-noise-ratio (SNR). The technique is then applied to directional emitters taking the imposed RSS modulation into account using a beampattern library. This results in significantly improved localization region confidence as compared to omnidirectional assumption approaches.